if 5th term of an AP is equal to 19 times the 19th term show that the 24th term of an AP is zero
Answer by mananth(16946)

if 5th term of an AP is equal to 19 times the 19th term show that the 24th term of an AP is zero
I think the problem should be
if 5 times the 5th term of an AP is equal to 19 times the 19th term show that the 24th term of an AP is zero
T5 = a+4d
t19= a+18d
5(a+4d)=19(a+18d)

5a +20d=19a+342d

14a+322d=0

14(a+23d)=0
a+23d=0
..
.

t24= a+23d
But a+23d=0
therefore
T24=0